OpenStudy (anonymous):

Graph y=^3 sqrt x-5-1

OpenStudy (jdoe0001):

\(\large {\textit{parent function}\implies \sqrt[3]{x}\\ \quad \\ \begin{array}{rllll} y=\sqrt[3]{x-5}&-1\\ \uparrow\ &\uparrow\\ \textit{horizontal shift}&\ \ \textit{vertical shift} \end{array}}\)

OpenStudy (anonymous):

What??? I cant read that?

OpenStudy (mertsj):

\[y=\sqrt[3]{x-5}-1\]

OpenStudy (mertsj):

Is that your problem?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[y=\sqrt[3]{x}\] Can you graph that?

OpenStudy (anonymous):

No I don't know how?

OpenStudy (mertsj):

Consider x=0, If x = 0 what is y?

OpenStudy (mertsj):

\[y=\sqrt[3]{0}\]

OpenStudy (anonymous):

Y=3?

OpenStudy (mertsj):

Do you know that extracting the third root and raising to the third power are inverse operations?

OpenStudy (anonymous):

Yes I know I just don't know how to graph it

OpenStudy (mertsj):

For example: \[\sqrt[3]{8}= 2 \because (2)^3=8\]

OpenStudy (mertsj):

So you don't know how to graph ordered pairs or what?